# Is Meta-training Really Necessary for Molecular Few-Shot Learning ?

### v1

@@ 3.3 Quadratic probing
$f_{2}\left({\mathbf{z}}_{i},\Theta\right)$ $=O(\sum_{k\in\{0,1\}}\lVert{\mathbf{M}}_{k}\rVert_{F}),$ (6)
$\text{and}\quad\tilde{f}_{2}\left({\mathbf{z}}_{i},\Theta\right)$ $=O(\sum_{k\in\{0,1\}}\log\lVert{\mathbf{M}}_{k}\rVert_{F}),$
@@
${\mathbf{M}}_{k}=\left(\frac{1}{\lvert\mathcal{S}_{y}\rvert}\sum_{i\in\mathcal{S}_{y}}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{k}\right)\left({\mathbf{z}}_{i}-{\mathbf{w}}_{k}\right)^{T}\right)^{-1}.$ (7)
@@
Minimizing $f_{1}+\tilde{f}_{2}$ is equivalent to maximizing the log-likelihood of a multivariate Gaussian distribution with mean ${\mathbf{w}}_{k}$ and covariance matrix $\Sigma_{k}$, explaining why the closed form solution we obtain corresponds to the estimator of the empirical covariance matrix’s inverse.
@@
Finally, we train the quadratic probe by updating ${\mathbf{M}}_{k}$ using the closed-form solution obtained above (Equation 7) and performing a gradient step on ${\mathbf{w}}_{k}$ using the cross-entropy loss. Since the estimation of the covariance matrix can be unstable, we use a shrinkage parameter $\lambda\in[0,1]$ to stabilize the estimation of the covariance matrix ($\Sigma_{k}\triangleq(1-\lambda)\Sigma_{k,\text{empirical}}+\lambda{\mathbf{I}}$). We consider $\lambda$ as a hyperparameter of the model to be optimized (see Appendix A.1 for further details).
@@
4 Experimental Results
4.1 FS-mol benchmark
Methods compared. We compare our methods to various meta-learning algorithms. First, we compare our approach to the baselines presented in the paper of Stanley et al. (2021). Their methods included: GNN-MAML (Finn et al., 2017), and Protonet (Snell et al., 2017). We considered the methods presented in the paper of Chen et al. (Chen et al., 2023), ADKT-IFT, the Property Aware Relation network (PAR(Wang et al., 2021), and conditional neural process (CNP) on this benchmark (Garnelo et al., 2018). Finally, we evaluated MHNFS (Schimunek et al., 2023) on this benchmark. We used the author’s scripts to train the model while implementing the evaluation scripts, which were unavailable on the author’s repository.
We report the results of the fine-tuning baselines already proposed by Stanley et al.(2021), such as fully fine-tuning of a molecular attention transformer (MAT(Maziarka et al., 2019) and GNN-MT, their multitask fine-tuning approach. We also re-evaluated the few-shot adaptation capabilities of CLAMP (Seidl et al., 2023) on the FS-mol benchmark (without adding the support set of each task to the model’s pre-training set).
We also consider simple baselines: the authors of FS-mol proposed to train a random forest classifier and a kNN on the support set’s molecular fingerprint   (Stanley et al., 2021). The authors also propose to train a full single-task GNN on the support set (GNN-ST). While proposing MHNFS, Schimuneck et al. (2023) evaluated the Similarity Search, a standard chemoinformatic relying only on the Tanimoto similarity between fingerprints (Bender et al., 2004; Cereto-Massagué et al., 2015).
Results Table 1 presents the results of the different methods on the FS-mol test set. First, our evaluation of CLAMP achieves results similar to the ones reported in the original paper with 16 labelled points and obtains better results with 64 and 128 labelled points. On the other hand, we did not succeed in reproducing MHNFS’s results with our implementation.
Figure 3: Mean performances of each method on the FS-mol benchmark, for support set sizes of 16, 32, 64 and 128.
@@ 4.1 FS-mol benchmark
As expected, the quadratic probe outperforms the linear probe as the support set size increases. This result is expected since the quadratic probe is a model fine-tuning more parameters than the linear probe. Additionally, the quadratic probe consistently reaches performances that are competitive with ADKT-IFT’s results at each support set size.
Figure 3 provides another visualization of the results, by using autorank (Herbold, 2020) to rank the methods based on their performances. This metric not only takes into account the mean performances of the methods, but also the consistency of the results across the different models, where we can see how fine-tuning methods achieve comparable results to meta-learning models while being more straightforward and more convenient in a black-box setting.
@@ 4.3 Impact of Domain Shifts
To further investigate the robustness of the different methods, we wish to evaluate the models’ performances when the tasks they are assessed on differ from their training set. We designed scenarios extracted from the therapeutic data commons platform (Huang et al., 2021) and the LIT-PCBA dataset (Tran-Nguyen et al., 2020), in which the tasks yield a domain shift compared to the base dataset.
First, the FS-mol dataset mainly comprises balanced tasks, as explained in the previous section, while the tasks we will consider in this assessment are more imbalanced. The robustness of those different methods to class imbalance is essential, as molecules are often inactive in practice. We also propose to assess the ability of these few-shot models to retrieve active compounds from a large molecular library (library screening) while having access to a few labelled examples. The ability to rank and identify the molecules based on their chances of being active can also be helpful while using generative models, generating large batches of molecules, to identify the most promising ones.
In this section, we focus our work on the top-performing models on FS-mol, the linear and quadratic probes (l-probe, q-probe), protonet, adkt-ift, clamp, and the standard chemo-informatic method: simsearch.
@@ 4.3.1 QSAR modelling with imbalanced class distribution
Results. The results of the different methods on those datasets are presented in Figure 5. While the state-of-the-art meta-learning methods achieve competitive results on the KIBA dataset, our baselines outperform all other methods when the distribution of the labels becomes more imbalanced. The performances of the linear and quadratic probes are similar, which is expected as we performed this evaluation on a support set size of 16, when both methods achieved similar results on the main benchmark. These results seem to indicate that the fine-tuning baselines are more robust to prior shifts than the meta-learning methods. This can be explained by the fact that the base dataset is divided into tasks during meta-training, giving the model information on the expected marginal distributions of positive and negative classes.
@@ 4.3.2 Library screening
Datasets. Finding potential hits (active molecules) in a large library of compounds, referred to as library screening, is a common step in drug discovery, where scientists can then buy the most promising compounds from a commercial molecular database (Gentile et al., 2022). In these tasks, the proportion of """"hits"""" (active compounds) is usually very low (1̃%), yielding a much more significant prior shift than in the previous section. Besides, on the opposite of the section above, the targets considered in the datasets we will use are not kinases (see Appendix A.2.2), including another type of domain shift. To perform this evaluation, we will use 6 HTS datasets from the therapeutic data commons platform (Huang et al., 2021) and ten datasets from the LIT-PCBA database (Tran-Nguyen et al., 2020); details on the task choice can be found in the Appendix A.2.2.
@@ 5 Summary and Concluding Remarks
While meta-learning methods consistently yield highly competitive results in molecular few-shot learning, it is essential to acknowledge that fine-tuning baselines have the potential to achieve comparable outcomes. In particular, we explored the intricacies of optimizing an extension of the linear probe, a quadratic probe. This extension is capable of competing with state-of-the-art meta-learning methods while relying on a model pre-trained with a standard multitask objective.
Moreover, our findings demonstrate that these fine-tuning baselines exhibit heightened robustness when confronted with a shift in the label distribution. This resilience further enhances their practical utility in diverse scenarios. Regrettably, these baseline approaches are often overlooked in the existing literature in this field. Notably, these methods possess distinctive advantages, particularly in a black-box setting, where they can be readily benchmarked upon the availability of a pre-trained model.
The anonymized codebase used in this paper is available at: https://github.com/Fransou/is-meta-learning-necessary
6 Impact Statements
This paper presents work that aims to advance the field of Machine Learning. Our work has many potential societal consequences. We hope our work can be used to help drug discovery pipelines, help discover treatments, and accelerate the process (which would limit the amount of animal testing, for instance).
@@ Appendix A Appendix
A.0.1 Proofs
@@
The samples from both classes are linearly separable. As such, there exists a vector ${\mathbf{v}}$ such that
$\forall i\in\{0,\ldots,N\},\begin{cases}{\mathbf{v}}^{T}{\mathbf{z}}_{i}\textgreater b&\text{ if }y_{i}=1\\ {\mathbf{v}}^{T}{\mathbf{z}}_{i}\textless b&\text{ if }y_{i}=0,\\ \end{cases}$ (8)
where $b$ is a constant, and we take ${\mathbf{v}}$ such that $\lVert{\mathbf{v}}\rVert=1$.
First, by dividing the expression of $p_{\Theta}(Y=y|Z={\mathbf{z}})$ by its numerator:
$\mathcal{L}_{ce}({\mathbf{z}}_{i},y_{i},{\Theta})=\log(1+e^{-(\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{1-y_{i}}\rVert^{2}_{{\mathbf{M}}_{1-y_{i}}}-\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{y_{i}}\rVert^{2}_{{\mathbf{M}}_{y_{i}}})})\geq 0.$ (9)
Our goal, is to build a set of parameters $\Theta(\lambda)$ such that $\lim_{\lambda\to+\infty}\mathcal{L}_{ce}({\mathbf{z}}_{i},y_{i},{\Theta(\lambda)})=0$, and $\lambda$ is an eigenvalue of ${\mathbf{M}}_{k}$ for all $k\in\{0,1\}$.
@@
We choose ${\mathbf{M}}_{k}(\lambda)$ so that its first eigenvector is ${\mathbf{v}}$ with an eigenvalue of $\lambda_{k,0}=\lambda$, and we keep all other eigenvalues and eigenvector constant. We set ${\mathbf{w}}_{k}$ as:
${\mathbf{w}}_{k}=\begin{cases}(b-1){\mathbf{v}}\quad\text{if }k=0\\ (b+1){\mathbf{v}}\quad\text{if }k=1.\\ \end{cases}$
@@
$\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}$ $=\frac{1}{2}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{k}\right)^{T}{\mathbf{U}}^{T}\Lambda_{k}{\mathbf{U}}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{k}\right)$
$\sim_{\lambda\to+\infty}\frac{\lambda}{2}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{k}\right)^{T}{\mathbf{v}}{\mathbf{v}}^{T}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{k}\right)$
$\sim_{\lambda\to+\infty}\frac{\lambda}{2}\left({\mathbf{v}}^{T}{\mathbf{z}}_{i}-{\mathbf{v}}^{T}{\mathbf{w}}_{k}\right)^{2}.$
@@
$\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{1-y_{i}}\rVert^{2}_{{\mathbf{M}}_{1-y_{i}}}$ $\sim_{\lambda\to+\infty}\frac{\lambda}{2}\left(1+{\mathbf{v}}^{T}{\mathbf{z}}_{i}-b\right)^{2}$
$\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{y_{i}}\rVert^{2}_{{\mathbf{M}}_{y_{i}}}$ $\sim_{\lambda\to+\infty}\frac{\lambda}{2}\left(1-\left({\mathbf{v}}^{T}{\mathbf{z}}_{i}-b\right)\right)^{2}.$
Since we chose $y_{i}=1$, the assumption of linear separability in the equation 8 states that ${\mathbf{v}}^{T}{\mathbf{z}}_{i}-b\textgreater 0$. As a result, $\left(1+{\mathbf{v}}^{T}{\mathbf{z}}_{i}-b\right)^{2}\textgreater\left(1-\left({\mathbf{v}}^{T}{\mathbf{z}}_{i}-b\right)\right)^{2}$. Hence, for all $i$ such that $y_{i}=1$,
$\left({\mathbf{v}}^{T}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{1-y_{i}}\right)\right)^{2}\textgreater\left({\mathbf{v}}^{T}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{y_{i}}\right)\right)^{2}.$ (10)
@@
$\lim_{\lambda\to+\infty}\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{1-y_{i}}\rVert^{2}_{{\mathbf{M}}_{1-y_{i}}}-\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{y_{i}}\rVert^{2}_{{\mathbf{M}}_{y_{i}}}$ $=\lim_{\lambda\to+\infty}\frac{\lambda}{2}\left(\left({\mathbf{v}}^{T}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{1-y_{i}}\right)\right)^{2}-\left({\mathbf{v}}^{T}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{y_{i}}\right)\right)^{2}\right)$ (11)
@@
$\lim_{\lambda\to+\infty}\mathcal{L}_{ce}({\mathbf{z}}_{i},y_{i},{\Theta(\lambda)})=0.$ (12)
@@
It is straightforward to see that for a given data point ${\mathbf{z}}_{i}$, $\mathcal{L}_{ce}({\mathbf{z}}_{i},y_{i},\Theta)=f_{1}\left({\mathbf{z}}_{i},y_{i},\Theta\right)+f_{2}\left({\mathbf{z}}_{i},\Theta\right)$.
@@
$0\leq\lVert z_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}=\lVert D_{k}\left(z_{i}-{\mathbf{w}}_{k}\right)\rVert_{2}^{2}\leq\lVert D_{k}\rVert_{F}^{2}\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert_{2}^{2}.$ (13)
@@
$\underset{{\mathbf{M}}_{k},k\in\{0,1\}}{\arg\min}f_{2}\left({\mathbf{z}}_{i},\Theta\right)$ $=\underset{{\mathbf{M}}_{k},k\in\{0,1\}}{\arg\min}\log\left(\sum_{k\in\{0,1\}}\exp\left(-\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)\right)$
$=\underset{{\mathbf{M}}_{k},k\in\{0,1\}}{\arg\min}\sum_{k\in\{0,1\}}\exp\left(-\lVert{\mathbf{z}}_{i}-{\textnormal{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)$
=$=\left\{\underset{{\mathbf{M}}_{k}}{\arg\min}\exp\left(-\lVert{\mathbf{z}}_{i}-{\textnormal{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)\right\}_{k\in\{0,1\}}.$
Minimizing $f_{2}$ is then equivalent to maximizing $\lVert{\mathbf{z}}_{i}-{\textnormal{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}$ for all $k\in\{0,1\}$, and:
$\lVert z_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}=({\mathbf{z}}_{i}-{\mathbf{w}}_{k})^{T}{\mathbf{M}}_{k}({\mathbf{z}}_{i}-{\mathbf{w}}_{k})\geq\sigma_{k,min}\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{2},$ (14)
@@
For a molecular embedding ${\mathbf{z}}_{i}$, the expression of $f_{2}\left({\mathbf{z}}_{i},\Theta\right)$ is:
$f_{2}\left({\mathbf{z}}_{i},\Theta\right)=\log\left(\sum_{k\in\{0,1\}}\exp\left(-\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)\right).$

### v2

@@ 3.3 Quadratic probing
$f_{2}\left({\bm{z}}_{i},\Theta\right)$ $=O(\sum_{k\in\{0,1\}}\lVert{\mathbf{M}}_{k}\rVert_{F}),$ (6)
$\text{and}\quad\tilde{f}_{2}\left({\bm{z}}_{i},\Theta\right)$ $=O(\sum_{k\in\{0,1\}}\log\lVert{\mathbf{M}}_{k}\rVert_{F}),$
@@
${\mathbf{M}}_{k}=\left(\frac{1}{\lvert\mathcal{S}_{y}|}\sum_{i\in\mathcal{S}_{y}}\left({\bm{z}}_{i}-{\bm{w}}_{k}\right)\left({\bm{z}}_{i}-{\bm{w}}_{k}\right)^{T}\right)^{-1}.$ (7)
@@
Minimizing $f_{1}+\tilde{f}_{2}$ is equivalent to maximizing the log-likelihood of a multivariate Gaussian distribution with mean ${\bm{w}}_{k}$ and covariance matrix $\Sigma_{k}$, explaining why the closed form solution we obtain corresponds to the estimator of the empirical covariance matrix’s inverse. Empirical results displaying the evolution of the largest eigenvalue of ${\mathbf{M}}_{k}$ using gradient descent and the quadratic probe are displayed in Appendix A.3
@@
Finally, we train the quadratic probe by updating ${\mathbf{M}}_{k}$ using the closed-form solution obtained above (Equation 7 and performing a gradient step on ${\bm{w}}_{k}$ using the cross-entropy loss. Since the estimation of the covariance matrix can be unstable, we use a shrinkage parameter $\lambda\in[0,1]$ to stabilize the estimation of the covariance matrix ($\Sigma_{k}\triangleq(1-\lambda)\Sigma_{k,\text{empirical}}+\lambda{\mathbf{I}}$). We consider $\lambda$ as a hyperparameter of the model to be optimized (refer to Appendix A.4 for further details). The algorithm of the training process of the quadratic probe can be found in Appendix A.1.
4 Experimental Results
4.1 FS-mol benchmark
Methods compared. We compare our methods to various meta-learning algorithms: GNN-MAML (Finn et al., 2017), Protonet (Snell et al., 2017), ADKT-IFT (Chen et al., 2023), the Property Aware Relation network (PAR) (Wang et al., 2021), and conditional neural process (CNP) (Garnelo et al., 2018), MHNFS (Schimunek et al., 2023). [^1] We used the author’s scripts to train the model while implementing the evaluation scripts, which were unavailable on the author’s repository. We report the results of fine-tuning baselines such as fully fine-tuning of a molecular attention transformer (MAT) (Maziarka et al., 2019) and GNN-MT, their multitask fine-tuning approach. We also re-evaluated the few-shot adaptation capabilities of CLAMP (Seidl et al., 2023) on the FS-mol benchmark (without adding the support set of each task to the model’s pre-training set). We also consider simple baselines: training a random forest classifier and a kNN on the support set’s molecular fingerprint (Stanley et al., 2021), and Similarity Search, a standard chemoinformatic method relying only on the Tanimoto similarity between these fingerprints (Bender et al., 2004; Cereto-Massagué et al., 2015).. The authors of FS-mol also propose to train a single-task GNN on the support set of a task as a baseline(GNN-ST).
@@ 4.1 FS-mol benchmark
Results. Table 1 presents the results of the different methods on the FS-mol test set. First, our evaluation of CLAMP achieves results similar to the ones reported in the original paper with 16 labelled points and obtains better results with 64 and 128 labelled points. On the other hand, we did not succeed in reproducing MHNFS’s results with our implementation.
@@
As expected, the quadratic probe obtains consistently slightly better results than the linear probe, and the difference becomes more significative as the amount of labeled data increases. This result is expected since the quadratic probe is a model fine-tuning more parameters than the linear probe, at the cost of a slightly more complex procedure (requiring to estimate ${\mathbf{M}}_{k}$). Additionally, the quadratic probe consistently reaches performances that are competitive with ADKT-IFT’s results at each support set size.
@@ 4.3 Impact of Domain Shifts
Table 2: Results on DTI tasks ($\Delta AUCPR$) for support set sizes of 16 and on the library screening tasks (ranks).
KIBA DAVIS BindingDB_Kd Library Screening
Model $|\mathcal{S}|=16$ $|\mathcal{S}|=16$ $|\mathcal{S}|=32$ $|\mathcal{S}|=64$ $|\mathcal{S}|=128$
Similarity search 17.3$\pm$ 0.9 19.1$\pm$ 3.1 17.3$\pm$ 1.3 3.97 3.69 3.19 2.78
Protonet NeurIPS2021 25.1$\pm$ 1.0 37.6$\pm$ 1.6 34.0$\pm$ 0.9 3.66 3.57 3.32 3.17
ADKF-IFT ICLR 2023 31.0$\pm$ 1.5 36.5$\pm$ 3.1 34.5$\pm$ 1.1 4.01 3.98 3.72 3.35
CLAMP ICLR 2023 26.1$\pm$ 1.3 37.0$\pm$ 2.0 34.2$\pm$ 1.2 3.42 3.19 3.28 3.42
Linear Probe 29.0$\pm$ 1.4 38.9$\pm$ 1.8 36.8$\pm$ 1.1 2.98 3.55 4.18 4.81
Quadratic Probe 28.8$\pm$ 1.4 39.2$\pm$ 1.8 37.1$\pm$ 1.0 2.96 3.02 3.31 3.46
To further investigate the robustness of the different methods, we wish to evaluate the models’ performances when the tasks they are evaluated on differ from the distribution of the training set. We designed scenarios extracted from the therapeutic data commons platform (Huang et al., 2021) and the LIT-PCBA dataset (Tran-Nguyen et al., 2020), in which the tasks yield a domain shift compared to the base dataset.
First, the FS-mol dataset mainly comprises balanced tasks, as explained in the previous section, while the tasks we will consider in this assessment are more imbalanced. The robustness of those different methods to class imbalance is essential, as molecules are often inactive in practice.
In this section, we focus our work on the top-performing models on FS-mol, the linear and quadratic probes (l-probe, q-probe), protonet, ADKF-IFT (adkt), clamp, and the standard chemo-informatic method: simsearch.
@@ 4.3.1 QSAR modelling with imbalanced class distribution
Results. The results of the different methods on those datasets are presented in Figure 3. While the state-of-the-art meta-learning methods achieve competitive results on the KIBA dataset, our baselines outperform all other methods when the distribution of the labels becomes more imbalanced. The performances of the linear and quadratic probes are similar, which is expected as we performed this evaluation on a support set consisting of 16 molecules, a setup where both methods achieved similar results on the main benchmark. These results seem to indicate that the fine-tuning baselines are more robust to prior shifts.
@@ 4.3.2 Library screening
Datasets. Finding potential hits (active molecules) in a large library of compounds, referred to as library screening, is a common step in drug discovery, where scientists can then buy the most promising compounds from a commercial molecular database (Gentile et al., 2022). In these tasks, the proportion of """"hits"""" (active compounds) is usually very low ($\approx 1\%$), yielding a much more significant prior shift than in the previous section. Besides, on the opposite of the section above, the targets considered in the datasets we will use are not kinases (see Appendix A.5.2), including another type of domain shift. To perform this evaluation, we will use 6 HTS datasets from the therapeutic data commons platform (Huang et al., 2021) and ten datasets from the LIT-PCBA database (Tran-Nguyen et al., 2020); details on the task choice can be found in the Appendix A.5.2.
@@ 5 Summary and Concluding Remarks
While meta-learning methods consistently yield highly competitive results in molecular few-shot learning, it is essential to acknowledge that fine-tuning baselines have the potential to achieve comparable outcomes. In particular, we explored the intricacies of optimizing an extension of the linear probe, a quadratic probe. Both probing methods are capable of competing with state-of-the-art meta-learning methods while relying on a model pre-trained with a standard multitask objective.
Moreover, our findings demonstrate that these fine-tuning baselines exhibit heightened robustness when confronted with a shift in the label distribution. This resilience further enhances their practical utility in diverse scenarios. Although the quadratic probe requires an additional step—estimating covariance matrices—it consistently yields slightly better results than the linear probe, with notable improvements under certain domain shifts. Thus, both methods provide efficient few-shot classifiers with distinct strengths that can be evaluated based on the demands of specific real-world applications.
Regrettably, these baseline approaches are often overlooked in the existing literature in this field. Notably, these methods possess distinctive advantages, particularly in a black-box setting, where they can be readily benchmarked upon the availability of a pre-trained model.
The anonymized codebase used in this paper is available at https://github.com/Fransou/Strong-Baseline-Molecular-FSL
6 Statement of Broader Impact
This paper presents work that aims to advance the field of Machine Learning. Our work has many potential societal consequences. We hope our work can be used to help drug discovery pipelines, help discover treatments, and accelerate the process. We did not identify any risk of harm specific to our work.
@@ Appendix A Appendix
A.1 Algorithm
Procedure 1 Training of the quadratic probe
0:  A support set $\mathcal{S}=\{\mathbf{z}_{i},y_{i}\}_{i\leq N}$, the model’s weights $\{\mathbf{w}_{k}\}_{k\in\{0,1\}}$, a learning rate $\alpha$ and a shrinkage coefficient $\lambda$.
  for $i\in\{1,\ldots,N_{epochs}\}$ do
     for $k\in\{0,1\}$  do
        $\mathcal{S}_{k}\leftarrow\{\mathbf{z}_{i}|y_{i}=k\}_{i\leq N}$
        ${\mathbf{M}}_{k}\leftarrow\left((1-\lambda)\frac{1}{\lvert\mathcal{S}_{k}|}\sum_{{\bm{z}}\in\mathcal{S}_{k}}\left({\bm{z}}-{\bm{w}}_{k}\right)\left({\bm{z}}-{\bm{w}}_{k}\right)^{T}+\lambda{\mathbf{I}}\right)^{-1}$Compute the covariance matrices}
     end for
     $\Theta\leftarrow\{{\bm{w}}_{k},{\mathbf{M}}_{k}\}_{k\in\{0,1\}}$
     $\mathcal{L}_{ce}(\Theta)\leftarrow\frac{1}{|\mathcal{S}|}\sum_{{\bm{z}},y\in\mathcal{S}}\mathcal{L}_{ce}({\bm{z}},y,{\Theta})$
     for $k\in\{0,1\}$ do
        ${\bm{w}}_{k}\leftarrow{\bm{w}}_{k}-\alpha\nabla_{w_{k}}\mathcal{L}_{ce}(\Theta)$Update the weights with gradient descent}
     end for
  end for
  Return $\Theta$
A.2 Proofs
@@
The samples from both classes are linearly separable. As such, there exists a vector ${\bm{v}}$ such that
$\forall i\in\{0,\ldots,N\},\begin{cases}{\bm{v}}^{T}{\bm{z}}_{i}\textgreater b&\text{ if }y_{i}=1\\ {\bm{v}}^{T}{\bm{z}}_{i}\textless b&\text{ if }y_{i}=0,\\ \end{cases}$ (8)
where $b$ is a constant, and we take ${\bm{v}}$ such that $\lVert{\bm{v}}\rVert=1$.
First, by dividing the expression of $p_{\Theta}(Y=y|Z={\bm{z}})$ by its numerator:
$\mathcal{L}_{ce}({\bm{z}}_{i},y_{i},{\Theta})=\log(1+e^{-(\lVert{\bm{z}}_{i}-{\bm{w}}_{1-y_{i}}\rVert^{2}_{{\mathbf{M}}_{1-y_{i}}}-\lVert{\bm{z}}_{i}-{\bm{w}}_{y_{i}}\rVert^{2}_{{\mathbf{M}}_{y_{i}}})})\geq 0.$ (9)
Our goal, is to build a set of parameters $\Theta(\lambda)$ such that $\lim_{\lambda\to+\infty}\mathcal{L}_{ce}({\bm{z}}_{i},y_{i},{\Theta(\lambda)})=0$, and $\lambda$ is an eigenvalue of ${\mathbf{M}}_{k}$ for all $k\in\{0,1\}$.
@@
We choose ${\mathbf{M}}_{k}(\lambda)$ so that its first eigenvector is ${\bm{v}}$ with an eigenvalue of $\lambda_{k,0}=\lambda$, and we keep all other eigenvalues and eigenvector constant. We set ${\bm{w}}_{k}$ as:
${\bm{w}}_{k}=\begin{cases}(b-1){\bm{v}}\quad\text{if }k=0\\ (b+1){\bm{v}}\quad\text{if }k=1.\\ \end{cases}$
@@
$\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}$ $=\frac{1}{2}\left({\bm{z}}_{i}-{\bm{w}}_{k}\right)^{T}{\mathbf{U}}^{T}\Lambda_{k}{\mathbf{U}}\left({\bm{z}}_{i}-{\bm{w}}_{k}\right)$
$\sim_{\lambda\to+\infty}\frac{\lambda}{2}\left({\bm{z}}_{i}-{\bm{w}}_{k}\right)^{T}{\bm{v}}{\bm{v}}^{T}\left({\bm{z}}_{i}-{\bm{w}}_{k}\right)$
$\sim_{\lambda\to+\infty}\frac{\lambda}{2}\left({\bm{v}}^{T}{\bm{z}}_{i}-{\bm{v}}^{T}{\bm{w}}_{k}\right)^{2}.$
@@
$\lVert{\bm{z}}_{i}-{\bm{w}}_{1-y_{i}}\rVert^{2}_{{\mathbf{M}}_{1-y_{i}}}$ $\sim_{\lambda\to+\infty}\frac{\lambda}{2}\left(1+{\bm{v}}^{T}{\bm{z}}_{i}-b\right)^{2}$
$\lVert{\bm{z}}_{i}-{\bm{w}}_{y_{i}}\rVert^{2}_{{\mathbf{M}}_{y_{i}}}$ $\sim_{\lambda\to+\infty}\frac{\lambda}{2}\left(1-\left({\bm{v}}^{T}{\bm{z}}_{i}-b\right)\right)^{2}.$
Since we chose $y_{i}=1$, the assumption of linear separability in the equation 8 states that ${\bm{v}}^{T}{\bm{z}}_{i}-b\textgreater 0$. As a result, $\left(1+{\bm{v}}^{T}{\bm{z}}_{i}-b\right)^{2}\textgreater\left(1-\left({\bm{v}}^{T}{\bm{z}}_{i}-b\right)\right)^{2}$. Hence, for all $i$ such that $y_{i}=1$,
$\left({\bm{v}}^{T}\left({\bm{z}}_{i}-{\bm{w}}_{1-y_{i}}\right)\right)^{2}\textgreater\left({\bm{v}}^{T}\left({\bm{z}}_{i}-{\bm{w}}_{y_{i}}\right)\right)^{2}.$ (10)
@@
$\lim_{\lambda\to+\infty}\lVert{\bm{z}}_{i}-{\bm{w}}_{1-y_{i}}\rVert^{2}_{{\mathbf{M}}_{1-y_{i}}}-\lVert{\bm{z}}_{i}-{\bm{w}}_{y_{i}}\rVert^{2}_{{\mathbf{M}}_{y_{i}}}$ $=\lim_{\lambda\to+\infty}\frac{\lambda}{2}\left(\left({\bm{v}}^{T}\left({\bm{z}}_{i}-{\bm{w}}_{1-y_{i}}\right)\right)^{2}-\left({\bm{v}}^{T}\left({\bm{z}}_{i}-{\bm{w}}_{y_{i}}\right)\right)^{2}\right)$ (11)
@@
$\lim_{\lambda\to+\infty}\mathcal{L}_{ce}({\bm{z}}_{i},y_{i},{\Theta(\lambda)})=0.$ (12)
@@
It is straightforward to see that for a given data point ${\bm{z}}_{i}$, $\mathcal{L}_{ce}({\bm{z}}_{i},y_{i},\Theta)=f_{1}\left({\bm{z}}_{i},y_{i},\Theta\right)+f_{2}\left({\bm{z}}_{i},\Theta\right)$.
@@
$0\leq\lVert z_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}=\lVert D_{k}\left(z_{i}-{\bm{w}}_{k}\right)\rVert_{2}^{2}\leq\lVert D_{k}\rVert_{F}^{2}\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert_{2}^{2}.$ (13)
@@
$\underset{{\mathbf{M}}_{k},k\in\{0,1\}}{\arg\min}f_{2}\left({\bm{z}}_{i},\Theta\right)$ $=\underset{{\mathbf{M}}_{k},k\in\{0,1\}}{\arg\min}\log\left(\sum_{k\in\{0,1\}}\exp\left(-\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)\right)$
$=\underset{{\mathbf{M}}_{k},k\in\{0,1\}}{\arg\min}\sum_{k\in\{0,1\}}\exp\left(-\lVert{\bm{z}}_{i}-{\textnormal{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)$
=$=\left\{\underset{{\mathbf{M}}_{k}}{\arg\min}\exp\left(-\lVert{\bm{z}}_{i}-{\textnormal{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)\right\}_{k\in\{0,1\}}.$
Minimizing $f_{2}$ is then equivalent to maximizing $\lVert{\bm{z}}_{i}-{\textnormal{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}$ for all $k\in\{0,1\}$, and:
$\lVert z_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}=({\bm{z}}_{i}-{\bm{w}}_{k})^{T}{\mathbf{M}}_{k}({\bm{z}}_{i}-{\bm{w}}_{k})\geq\sigma_{k,min}\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{2},$ (14)
@@
For a molecular embedding ${\bm{z}}_{i}$, the expression of $f_{2}\left({\bm{z}}_{i},\Theta\right)$ is:
$f_{2}\left({\bm{z}}_{i},\Theta\right)=\log\left(\sum_{k\in\{0,1\}}\exp\left(-\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)\right).$